OpenStudy (anonymous):

You started to plan for your retirement 2 years ago by investing into a individual retirement account (IRA). You deposited $2,520 dollars over the two year period and your money grew to $2,850 over that same period. Determine, by what percent, did you money grow by for the two year period. Round your answer to the nearest percent

OpenStudy (anonymous):

will give medals

OpenStudy (anonymous):

a. 1.13% c. 13% b. 1.3% d. 130%

OpenStudy (anonymous):

@swagmaster47

OpenStudy (anonymous):

ILL FAN YOU

OpenStudy (anonymous):

@Harsha19111999

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (harsha19111999):

Hey

OpenStudy (harsha19111999):

You have got interest on your money

OpenStudy (anonymous):

hey wait

OpenStudy (harsha19111999):

You need by how my your money has raised

OpenStudy (harsha19111999):

I properly read the question just now

OpenStudy (anonymous):

need that formula back

OpenStudy (harsha19111999):

So change in money = (2850-2520/2520)* 100 = 13 percent

OpenStudy (anonymous):

can i jst have the answer

OpenStudy (harsha19111999):

ie. Option C. BTW the formula is I = PTR/100

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

yeah ima need the answer cause im lost

OpenStudy (anonymous):

@harsha19111999

OpenStudy (harsha19111999):

Its Option c